# Efficiency Meets Fidelity: A Novel Quantization Framework for Stable Diffusion

### v1

@@ Abstract
1 Introduction
@@
Diffusion models have yielded remarkable achievements and demonstrated exceptional performance across various generative tasks, [5, 14, 40, 42, 46, 47], particularly in the realm of text-to-image generation [5, 40, 42]. Nonetheless, these models often entail significant computational expenses, primarily due to two factors. Firstly, within a Diffusion model, a UNet [7, 41] carries out a time-consuming iterative sampling process to progressively denoise a random latent variable. Secondly, the pursuit of superior image quality and higher resolutions has resulted in larger model sizes, necessitating extensive time and memory resources. These challenges render Diffusion models (e.g., Stable Diffusion [40] and Stable Diffusion XL [38]) computationally demanding and difficult to deploy in real-world applications requiring low latency and scalability.
@@
Recently, many researchers have investigated quantization strategies for compressing Diffusion models [12, 24, 44, 50, 48, 11], predominantly utilizing Post-Training Quantization (PTQ) [25, 34]. PTQ does not require retraining or fine-tuning the network and therefore is more attractive than Quantization-Aware Training (QAT) [36] for large models. However, PTQ methods experience substantial performance degradation at 4 bits and below. Furthermore, the quantization of large text-to-image models, such as Stable Diffusion XL 1.0, can still require 1 day.
Meanwhile, another issue is that most existing research primarily concentrates on optimizing quantized models for high-quality image generation, paying little attention to the consistency of results produced by quantized and floating-point models. In the context of content creation and design, ensuring consistency in expected results is of paramount importance. It is imperative that quantized models exhibit a high degree of similarity to the style and content of images generated by floating-point models. Otherwise, users will encounter significant challenges in predicting and controlling the final results, necessitating extensive debugging and modification of cues, which will inevitably impact their productivity and creative expression. Moreover, alterations in the style of images generated by a quantized model will affect the performance of the downstream tasks [59, 56] and the overall reliability of the system.
To address the aforementioned issues, we propose a novel Stable Diffusion quantization framework that is specifically designed to achieve high fidelity and efficiency. We analyze the strengths and weaknesses of the existing pipeline for the joint optimization of the UNet. By leveraging their advantages, we propose our Serial-to-Parallel pipeline, which ensures consistency in the generated outputs while improving training stability. To further enhance fidelity, several techniques are introduced, including the preservation of temporal information, the utilization of multiple time-step activation quantizers, and a Hessian-based mixed-precision strategy.
@@
The quality of the generated results is evaluated in terms of both distributional and visual similarity. In comparison to previous PTQ methods, our framework demonstrates superior generation consistency in shorter training times across multiple Stable Diffusion models.
2 Related Work
2.1 Diffusion Model Acceleration
@@
The first method involves designing advanced samplers for pre-trained models, such as analytical trajectory estimation [2, 1], implicit sampler [46, 22, 61, 53], stochastic differential equations [47, 17, 19] and ordinary differential equations [30, 28, 60]. Although these methods can reduce the number of sampling iterations required, the significant parameter count and computational demands of Stable Diffusion models limit their application on edge devices.
@@
The second method involves retraining the model, such as diffusion scheme optimization [6, 8, 63, 33], knowledge distillation [31, 43], sample trajectory optimization [53, 23], and noise scale adjustment [21, 37]. Though these techniques effectively speed up the sampling process, re-training a Diffusion model is computationally intensive, especially for resource-constrained devices.
2.2 Diffusion Model quantization
@@
Quantization is a widely used technique that aids in reducing memory usage and speeding up computation. It is generally categorized into two types: QAT [9, 29, 16, 64, 58] and PTQ [25, 34, 15, 54, 27]. EfficientDM [11] is representative of QAT work, it proposes a data-free distillation framework and applies a quantization-aware variant of the low-rank adapter. While QAT is time-consuming and computationally heavy, recent studies focus on PTQ for Diffusion models, which does not require fine-tuning and only necessitates a small amount of unlabeled data for calibration. PTQ4DM [44] and Q-diffusion [24] focus on sampling the noise of the floating-point model across different timesteps, Q-diffusion further propose to split the activation of shortcut layers. PTQD [12] disentangle the quantization noise into its correlated and residual uncorrelated parts and correct them individually. PCR [48] progressively calibrates the activation quantizer considering the accumulated quantization error across timesteps and selectively relaxing the bit-width for several of those timesteps. However, these PTQ methods seldom consider the consistency of the generated output, and many of them are not designed for new, large pre-trained text-to-image models, such as Stable Diffusion.
Figure 1: Overview of our proposed quantization framework. (a) Dataset Generation: During the inference of the floating-point model, latent generated from various timesteps for each prompt are randomly sampled. (b) Time information precalculation: The feature map of time projection layers is precalculated for training and inference. Subsequently, the time embedding and projection layers are removed from UNet. (c) Serial-to-Parallel training pipeline: A Hessian-based sensitive ranking is assigned to each layer as well as different bit-widths. At each iteration, latent from various timesteps along with the corresponding prompts are selected from the dataset. The Loss function is calculated between the output and the sensitive layers.
3 Preliminaries
3.1 Diffusion models
@@
Diffusion models [14, 46] gradually add Gaussian noise with a variance schedule $\beta_{1},\ldots,\beta_{T}\in$ $(0,1)$ to real image $x_{0}\sim q(x)$ for $T$ times as sampling process, resulting in a sequence of noisy samples $x_{1},\ldots,x_{T}$. In DDPMs [14], the sampling process is a Markov chain, which can be formulated as:
$q\left(x_{1:T}\mid x_{0}\right)$ $=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right),q\left(x_{t}\mid x_{t-1}\right)$ (1)
@@
$\boldsymbol{\mu}_{\theta}\left(x_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(x_{t},t\right)\right)$ (2)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ is a trainable model to predict noise. The variance $\Sigma_{\theta}\left(\mathrm{x}_{t},t\right)$ can be either learned [37] or fixed to a constant schedule [14] $\sigma_{t}$. When it uses a constant schedule, $x_{t-1}$ can be expressed as:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(x_{t},t\right)\right)+\sigma_{t}\mathbf{z}$ (3)
@@
The formulas outlined in our research are based on the DDPM framework but can be easily adjusted for other accelerated sampling techniques such as DDIM [46], PNDM [28], and Euler [18].
3.2 Model Quantization
Quantization [35] is a key technique in model compression. This method compresses neural networks by reducing the number of bits used for model weights and activations. The quantization process can be formulated as:
$w_{q}=\mathrm{clip}\left(\mathrm{round}\left({\frac{w}{s}}\right)+z,q_{\mathrm{min}},q_{\mathrm{max}}\right)$ (4)
@@ 3.2 Model Quantization
$\hat{w}=(w_{q}-z)\times s$ (5)
We utilize uniform quantization in all our study experiments.
4 Method
@@
As illustrated in Fig. 1, we present a novel quantization framework for large pre-trained text-to-image diffusion models, including Stable Diffusion v1-4 and Stable Diffusion XL. We begin by introducing a Serial-to-Parallel training pipeline that not only addresses the consistency between the training and inference processes but also guarantees stability during training. Subsequently, several techniques are integrated into the pipeline. Multi-timestep activation quantizer is set to separately optimize the parameters associated with each timestep. Additionally, the time feature is precalculated and the accurate projection information is saved for training and inference. Furthermore, we implement a mixed-precision quantization strategy that assigns higher bit-width to sensitive layers and lower bit-width to insensitive layers.
4.1 Serial-to-Parallel Training Pipeline
Figure 2: Comparision of ’Serial’ and ’Parallel’ training pipeline.
Previous works on jointly optimizing quantized models and distillation-based compression can be roughly divided into two categories: (a) ’Serial’(e.g. [11]) and (b) ’Parallel’(e.g. [20, 32]), as illustrated in Fig. 2. The serial pipeline operates in a data-free manner,, requiring only a few prompts to generate the latent of the floating-point model. This latent is then used as input for the quantized model, which is updated in a chronological sequence. In contrast, the parallel pipeline is more closely aligned with the original Stable Diffusion training process. This approach relies on an image-text pair dataset, where the image is processed through a Variational Autoencoder (VAE) to derive the initial latent. In each iteration, multiple timesteps are randomly sampled, and the latent is then augmented with varying levels of Gaussian noise, as determined by the scheduler.
Figure 3: Difference in input latent range at each timestep with the same initial latent. (a) Gradually adding Gaussian noise based on Eq. 1. (b) Step-by-step denoising during inference of floating-point Stable Diffusion v1-4.
@@ 4.1 Serial-to-Parallel Training Pipeline
Figure 4: Box plot illustrating the gradient variations of the ’down_blocks.0.attentions.0.proj_in’ layer in the quantized Stable Diffusion v1-4 model during training.(a) represents the serial pipeline, and (b) represents the parallel pipeline.
Both frameworks have their own advantages and disadvantages. As illustrated in Fig. 3, the theoretical latent generated by adding Gaussian noise differs markedly from the actual latent range when reasoning with the floating-point model.
$\boldsymbol{\mu}_{fp}\left(x_{t},t\right)\neq x_{t-1}\leftarrow\mathcal{N}\left(x_{t};\sqrt{\alpha_{t}}x_{t-1},\beta_{t}\mathbf{I}\right)$ (6)
Consequently, it is more beneficial to use the latent from the floating-point model as input during the distillation process. This approach enhances the consistency of the quantized model’s outputs in comparison to those of the floating-point model.
@@
Also, in Stable Diffusion models, where all timesteps share the same weight, it is more appropriate to average the gradients across multiple timesteps rather than relying solely on their sequential order. As demonstrated in Fig. 4, we have documented the changes in gradients for both serial and parallel pipelines. It can be observed that the gradients remain relatively stable during parallel training, whereas they exhibit periodic oscillations during serial training. Previous research [39, 55] have indicated that Adam optimizer may sometimes sometimes underperform in the presence of periodic oscillating gradients.
@@
It can thus be concluded that the latent of the serial pipeline is more appropriate while the training procedure of the parallel pipeline is more reasonable. Building on this analysis, we introduce our method, termed ’Serial-to-Parallel’, which harnesses the strengths of both serial and parallel pipelines. The advantages of different pipelines are summarized in Tab. 1. Initially, the inference is conducted with the floating-point model, whereby the latent is randomly sampled from various timesteps for each prompt. During the training process, at each iteration, the latent is sampled from different timesteps along with their corresponding prompts from the latent dataset. This strategy renders our framework data-free, relying solely on prompts, while simultaneously enhancing generation consistency and ensuring training stability.
@@
Serial X
Parallel X X
Ours
Table 1: Comparison of our pipeline and previous pipeline.
4.2 Components For Higher Fidelity
Moreover, a variety of techniques are employed to guarantee the fidelity of the generated results.
Accurate activation quantization. Previous studies on Diffusion models [44, 24, 45, 57, 51] have shown that the activation distribution at different timesteps varies greatly, posing a challenge for activation quantization. We adopt different activation quantization parameter sets for different timesteps, which can be expressed as:
$\mathbf{s}_{l}=\left\{s_{l}^{0},s_{l}^{1},\ldots,s_{l}^{T-1}\right\},\mathbf{z}_{l}=\left\{z_{l}^{0},z_{l}^{1},\ldots,z_{l}^{T-1}\right\}$ (7)

### v2

@@ Abstract.
Abstract.
Text-to-image generation via Stable Diffusion models (SDM) have demonstrated remarkable capabilities. However, their computational intensity, particularly in the iterative denoising process, hinders real-time deployment in latency-sensitive applications. While Recent studies have explored post-training quantization (PTQ) and quantization-aware training (QAT) methods to compress Diffusion models, existing methods often overlook the consistency between results generated by quantized models and those from floating-point models. This consistency is paramount for professional applications where both efficiency and output reliability are essential. To ensure that quantized SDM generates high-quality and consistent images, we propose an efficient quantization framework for SDM. Our framework introduces a Serial-to-Parallel pipeline that simultaneously maintains training-inference consistency and ensures optimization stability. Building upon this foundation, we further develop several techniques including multi-timestep activation quantization, time information precalculation, inter-layer distillation, and selective freezing, to achieve high-fidelity generation in comparison to floating-point models while maintaining quantization efficiency.
Through comprehensive evaluation across multiple Stable Diffusion variants (v1-4, v2-1, XL 1.0, and v3), our method demonstrates superior performance over state-of-the-art approaches with shorter training times. Under W4A8 quantization settings, we achieve significant improvements in both distribution similarity and visual fidelity, while preserving a high image quality.
Diffusion model, quantization, image generation, edge hardware.
ccs: Computing methodologies Computer visionccs: Computer systems organization Embedded and cyber-physical systems
1. Introduction
@@
Diffusion models have yielded remarkable achievements and demonstrated exceptional performance across various generative tasks, (Chen et al., 2024; Ho et al., 2020; Rombach et al., 2022; Ruiz et al., 2023; Song et al., 2022, 2021), particularly in the realm of text-to-image generation (Chen et al., 2024; Rombach et al., 2022; Ruiz et al., 2023). Nonetheless, these models often entail significant computational expenses, primarily due to two factors. Firstly, within a Diffusion model, a UNet (Dhariwal and Nichol, 2021; Ronneberger et al., 2015) carries out a time-consuming iterative sampling process to progressively denoise a random latent variable. Secondly, the pursuit of superior image quality and higher resolutions has resulted in larger model sizes, necessitating extensive time and memory resources. These challenges render Diffusion models (e.g., Stable Diffusion (Rombach et al., 2022) and Stable Diffusion XL (Podell et al., 2023)) computationally demanding and difficult to deploy in real-world applications requiring low latency and scalability.
@@
Recently, many researchers have investigated quantization strategies for compressing Diffusion models (He et al., 2024; Li et al., 2023; Shang et al., 2023; Wang et al., 2024b; Tang et al., 2024; He et al., 2023; Sui et al., 2024), predominantly utilizing Post-Training Quantization (PTQ) (Li et al., 2021; Nagel et al., 2020). PTQ does not require a whole dataset and heavy retraining, making it more appealing than Quantization-Aware Training (QAT) (Nagel et al., 2022) for large models. However, PTQ methods on diffusion models experience substantial performance degradation at 4 bits and below. And their calibration processes remain computationally intensive building upon techniques like AdaRound (Nagel et al., 2020) and BRECQ (Li et al., 2021). The calibration process of large text-to-image models such as Stable Diffusion XL 1.0 still requires 1 day due to the necessary block-wise training procedures.
Meanwhile, most existing research focuses on optimizing quantized models for high-quality image generation, often neglecting the consistency between quantized and floating-point outputs. In content creation and design, such consistency is critical—quantized models must closely match the style and content of their floating-point counterparts. Otherwise, users face difficulties in predicting and controlling results, requiring extensive prompt tuning and adjustments, which hampers productivity and creative workflow. Moreover, discrepancies in generated styles can degrade the performance of downstream tasks (Zhang et al., 2023b; Wu et al., 2024) and undermine system reliability.
To overcome these limitations, we propose a novel Stable Diffusion quantization framework that is specifically designed to achieve high fidelity and efficiency. Through systematic analysis of existing pipelines for the joint optimization of quantized diffusion models, we identify two critical barriers that hinder high-fidelity quantization: ❶ Timestep-sequential fine-tuning leads to periodic oscillations in weight gradients, which in turn results in substantial instability in quantized weights. ❷ Distributional discrepancy between noise-scheduled training latents and inference-phase latents leads to activation misalignment. Addressing the revealed challenges, we propose our Serial-to-Parallel pipeline, which maintains training-inference consistency and ensures optimization stability. To further enhance fidelity, several techniques are integrated into the pipeline, including the preservation of temporal information, the utilization of multiple time-step activation quantizers, inter-layer distillation on sensitive layers, and a selective weight freezing strategy on these layers.
@@
The generated results are evaluated in terms of similarity, quality, and text-image alignment. In comparison to previous PTQ methods, our framework demonstrates superior generation consistency in shorter training times across UNet-based and MMDiT-based Stable Diffusion models.
Our contributions are summarized as follows:
  • We propose a Serial-to-Parallel pipeline that maintains training-inference consistency and ensures optimization stability.
  • We integrate several techniques into the pipeline for higher generation fidelity.
  • We validate the fidelity and consistency improvements over sota methods.
2. Related Work
2.1. Diffusion Model Acceleration
@@
The first method involves designing advanced samplers for pre-trained models, such as analytical trajectory estimation (Bao et al., 2022b, a), implicit sampler (Song et al., 2022; Kong and Ping, 2021; Zhang et al., 2023c; Watson et al., 2021), stochastic differential equations (Song et al., 2021; Jolicoeur-Martineau et al., 2021; Kim and Ye, 2022) and ordinary differential equations (Lu et al., 2022; Liu et al., 2022; Zhang and Chen, 2023). Although these methods can reduce the number of sampling iterations required, the significant parameter count and computational demands of Stable Diffusion models limit their application on edge devices.
@@
The second method involves retraining the model, such as diffusion scheme optimization (Chung et al., 2022; Franzese et al., 2023; Zheng et al., 2023; Lyu et al., 2022), knowledge distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022), sample trajectory optimization (Watson et al., 2021; Lam et al., 2022), and noise scale adjustment (Kingma et al., 2021; Nichol and Dhariwal, 2021). Though these techniques effectively speed up the sampling process, re-training a Diffusion model is computationally intensive, especially for resource-constrained devices.
2.2. Diffusion Model Quantization
@@
Quantization is a widely used technique that aids in reducing memory usage and speeding up computation. It is generally categorized into two types: QAT (Gong et al., 2019; Louizos et al., 2018; Jacob et al., 2018; Zhuang et al., 2018; Zhang et al., 2023a) and PTQ (Li et al., 2021; Nagel et al., 2020; Hubara et al., 2020; Wei et al., 2023; Lin et al., 2023). EfficientDM (He et al., 2023) is representative of QAT work, it proposes a data-free distillation framework and applies a quantization-aware variant of the low-rank adapter. While training on the whole dataset is time-consuming and computationally heavy, recent studies focus on PTQ for Diffusion models, which only necessitates a small amount of unlabeled data for calibration. PTQ4DM (Shang et al., 2023) and Q-diffusion (Li et al., 2023) focus on sampling the noise of the floating-point model across different timesteps, Q-diffusion further propose to split the activation of shortcut layers. PTQD (He et al., 2024) disentangle the quantization noise into its correlated and residual uncorrelated parts and correct them individually. PCR (Tang et al., 2024) progressively calibrates the activation quantizer considering the accumulated quantization error across timesteps and selectively relaxing the bit-width for several of those timesteps. However, these PTQ methods seldom consider the consistency of the generated output, and many of them are not designed for new, large pre-trained text-to-image models, such as Stable Diffusion.
3. Preliminaries
3.1. Diffusion Models
@@
Diffusion models (Ho et al., 2020; Song et al., 2022) gradually add Gaussian noise with a variance schedule $\beta_{1},\ldots,\beta_{T}\in$ $(0,1)$ to real image $x_{0}\sim q(x)$ for $T$ times as sampling process, resulting in a sequence of noisy samples $x_{1},\ldots,x_{T}$. In DDPMs (Ho et al., 2020), the sampling process is a Markov chain, which can be formulated as:
(1) $q\left(x_{1:T}\mid x_{0}\right)$ $=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right),q\left(x_{t}\mid x_{t-1}\right)$
@@
(2) $\boldsymbol{\mu}_{\theta}\left(x_{t},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(x_{t},t\right)\right)$
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ is a trainable model to predict noise. The variance $\Sigma_{\theta}\left(\mathrm{x}_{t},t\right)$ can be either learned (Nichol and Dhariwal, 2021) or fixed to a constant schedule (Ho et al., 2020) $\sigma_{t}$. When it uses a constant schedule, $x_{t-1}$ can be expressed as:
(3) $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(x_{t},t\right)\right)+\sigma_{t}\mathbf{z}$
@@
The formulas outlined in our research are based on the DDPM framework but can be easily adjusted for other accelerated sampling techniques such as DDIM (Song et al., 2022), PNDM (Liu et al., 2022), and Euler (Karras et al., 2022).
In this work, we adopt Denoising Diffusion Block (DDB) as the unified term for the backbone denoising networks $\boldsymbol{\epsilon}_{\theta}(\cdot)$. Specifically, for SD1.4, SD2.1, and SDXL, the DDB is implemented as a U-Net architecture. For SD3, the DDB follows the MM-DiT (Multi-Modal Diffusion in Transformer) design.
3.2. Model Quantization
Quantization (Nagel et al., 2021) is a key technique in model compression. This method compresses neural networks by reducing the number of bits used for model weights and activations. The quantization process can be formulated as:
(4) $w_{q}=\mathrm{clip}\left(\mathrm{round}\left({\frac{w}{s}}\right)+z,q_{\mathrm{min}},q_{\mathrm{max}}\right)$
@@ 3.2. Model Quantization
(5) $\hat{w}=(w_{q}-z)\times s$
4. Methodology
@@
We begin by analyzing the advantages and disadvantages of previous training pipelines for large pre-trained text-to-image diffusion models. Then we introduce our Serial-to-Parallel training pipeline which combines their advantages. Subsequently, several techniques are integrated into the pipeline. Multi-timestep activation quantizer is set to separately optimize the parameters associated with each timestep. Additionally, the time feature is precalculated and the accurate projection information is saved for training and inference. Furthermore, feature distillation on sensitive layers is employed and selective freezing is applied to those sensitive layers.
4.1. Serial-to-Parallel Pipeline
Figure 1. Comparison of (A) ’Serial’ pipeline which mimics the inference process of SDM and (B) ’Parallel’ pipeline which is more aligned with the pretraining process of SDM. $X_{t}^{n}$ denotes the $n^{th}$ noisy input at timestep $t$. $P$ denotes prompts. $\mu$ denotes the predicted noise.
@@ 4.1. Serial-to-Parallel Pipeline
Figure 2. Box plot illustrating the gradient variations in the quantized Stable Diffusion v1-4 model during training. (A) represents the serial pipeline, and (B) represents the parallel pipeline.
Previous works on jointly optimizing quantized models and distillation-based compression can be roughly divided into two categories: (a) ’Serial’(e.g. (He et al., 2023)) and (b) ’Parallel’(e.g. (Kim et al., 2023; Luo et al., 2023)), as illustrated in Fig. 1. The serial pipeline mimics the inference of Stable Diffusion models, and operates in a data-free manner. As shown in Fig. 1. (A), the initial latent $\{X_{0}^{1}...X_{0}^{n}\}$ are random noises, and only a few prompts are required as input. The latent is denoised step by step based on the prediction of full-precision models, which is sent into quantized DDB. In contrast, the parallel pipeline is more closely aligned with the original Stable Diffusion training process. As shown in Fig. 1. (B), this approach relies on an image-text pair dataset, where the image is processed through a Variational Autoencoder (VAE) to derive the initial latent. In each iteration, multiple timesteps are randomly sampled, and the latent is then augmented with varying levels of Gaussian noise, as determined by the scheduler. We will present two key observations to reveal their own advantages and disadvantages.
Figure 3. Comparison of loss (left) and weight oscillation (right) between serial and parallel pipeline. Serial pipeline suffers from severe oscillation
@@
Observation ❶: In serial pipeline, periodic oscillations of gradients leads to severe weight oscillations. As demonstrated in Fig. 2, we have documented the changes in gradients for both serial and parallel pipelines. It can be observed that the gradients remain relatively stable during parallel training, whereas they exhibit periodic oscillations during serial training. Previous research (Reddi et al., 2019; Wilson et al., 2017) have indicated that the Adam optimizer may underperform in the presence of periodic oscillating gradients.
Building upon Osqat (Nagel et al., 2022), we systematically analyze weight oscillation patterns across 30 layers, as shown in Fig. 3 (right). The serial pipeline demonstrates progressively accumulating weight oscillations, ultimately retaining 9% unstable parameters - a phenomenon directly correlated with its deteriorating training stability. As evidenced in the green highlighted region of Fig. 3 (left), this manifests as persistent loss fluctuations beginning at mid-training that persistently hinder stable convergence. In contrast, the parallel pipeline averages the gradients across multiple time-steps at each iteration and achieves better optimization stability, exhibiting both significantly reduced oscillation rates and smoother loss descent trajectories.
Figure 4. Difference in noisy input range at each timestep with the same initial latent. (a) Adding Gaussian noise based on Eq. 1. (b) Step-by-step denoising during inference.
Figure 5. Overview of our quantization framework. (A) Serial dataset generation: During the inference of floating-point models, latent generated from various timesteps for each prompt are randomly sampled. (B) Time information precalculation: The feature map of time projection layers is precalculated for training and inference. (C) Parallel training: At each iteration, latent from various timesteps along with the corresponding prompts are selected from the dataset. The Loss function is calculated between the output and the sensitive layers. Iterative freezing is applied to these sensitive layers for better stability.
Observation ❷: In parallel pipeline, the input latent range differs between training and inference. During training, noisy inputs are generated by adding Gaussian noise to image-derived latents, whereas inference employs randomly initialized latents undergoing progressive denoising. As shown in Fig. 4, these two approaches produce significantly different numerical ranges in the latent representations, that is:
(6) $\boldsymbol{\mu}_{fp}\left(x_{t},t\right)\neq x_{t-1}\leftarrow\mathcal{N}\left(x_{t};\sqrt{\alpha_{t}}x_{t-1},\beta_{t}\mathbf{I}\right)$
This mismatch poses challenges for activation quantization, as the quantizer parameters, calibrated on training data, could become suboptimal for inference distributions, potentially amplifying quantization errors. Notably, this issue does not affect the serial pipeline, where training and inference latent ranges remain consistent. Consequently, it is more beneficial to use the latent sampled from the floating-point model inference as noisy input.
@@
It can thus be concluded that the input latent of the serial pipeline is more appropriate while the training procedure of the parallel pipeline is more reasonable. To synergize these advantages, we propose our ’Serial-to-Parallel’ pipeline that (1) Generates latents serially. As shown in Fig. 5 (A), the inference is conducted with the floating-point model, whereby the latent is randomly sampled from various timesteps for each prompt. (2) Trains in parallel by utilizing these pre-computed latents. As shown in Fig. 5 (C), at each iteration, the latent is sampled from different timesteps along with their corresponding prompts from the latent dataset. This strategy renders our framework data-free, relying solely on prompts, while simultaneously enhancing generation consistency and ensuring training stability. The comparison of different pipelines is summarized in Tab. 1. Our serial-to-parallel pipeline effectively enhances generation consistency while remaining data-free.
Table 1. Comparison of our pipeline and previous pipeline on Stable Diffusion v2.1 with w4a8 quantization. Osc.: percentage of oscillating weights at the end of training. Please refer to section5.1 for details of evaluation metrics.
@@
[…]
Serial 13.68 0.50 0.46 9.05
Parallel X 12.92 0.53 0.42 0.33
Serial-to -Parallel 12.05 0.55 0.40 0.29
4.2. Components For Higher Fidelity
Moreover, a variety of techniques are employed to further enhance the fidelity of the generated results.
Accurate Activation Quantization. Previous studies on Diffusion models (Shang et al., 2023; Li et al., 2023; So et al., 2024; Yang et al., 2023; Wang et al., 2024a) have shown that the activation distribution at different timesteps varies greatly, posing a challenge for activation quantization. We adopt different activation quantization parameter sets for different timesteps, which can be expressed as:
(7) $\mathbf{s}_{l}=\left\{s_{l}^{0},s_{l}^{1},\ldots,s_{l}^{T-1}\right\},\mathbf{z}_{l}=\left\{z_{l}^{0},z_{l}^{1},\ldots,z_{l}^{T-1}\right\}$